Opened 10 years ago

Closed 9 years ago

#403 closed flight processing (fixed)

GB03/01, flight day 165a/2011, Fiskerton Barlings

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors: emca

Description (last modified by emca)

Data location: ~arsf/arsf_data/2011/flight_data/uk/GB03_01-2011_165a_Fiskerton_Barlings

Data arrived from ARSF via network transfer on 15/06/2011

GB03/01 Scientific objective:
? (no 2003 application forms)

GB03/01 Priority: a4h

PI: Jim Williams

Sensors:

Camera (11/07/2011)
Eagle (11/07/2011)
Hawk (11/07/2011)
LiDAR (22/09/2011)

Attachments (1)

2011165.sync (521 bytes) - added by iopa 10 years ago.
Eagle & Hawk SCT offsets

Download all attachments as: .zip

Change History (20)

comment:1 Changed 10 years ago by knpa

  • Description modified (diff)

comment:2 Changed 10 years ago by emca

PI has requested the LiDAR data as LAS files in WGS84 projection. He would also like the SBET file to be included with the delivery.

comment:3 Changed 10 years ago by emca

Processed IPAS data and created sol, sbet and camera sol file.

comment:4 Changed 10 years ago by iopa

Renamed RAW hyperspectral files to include prefixes 'VNIR' and 'SWIR' respectively for eagle/hawk.

Changed 10 years ago by iopa

Eagle & Hawk SCT offsets

comment:5 Changed 10 years ago by iopa

Removed pre-flight raw images;
1-22
Removed raw images over cloud or saturated;
198,249,253-254,290-291,347,363

Camera ready for checking.

comment:6 Changed 10 years ago by anch

Camera has been checked.

No problems with the delivery, so ready to deliver

comment:7 Changed 10 years ago by knpa

  • Owner set to iopa
  • Status changed from new to assigned

Finished HS delivery check.

All looks good apart from the following:

  1. Yellow tint on one side of hawk in the screenshots
  2. Blank first and last band for eagle lev1s
  3. First band hawk blank lev1

comment:8 Changed 10 years ago by iopa

The first/last bands referred to are masked out according to new calibration of the sensors.

comment:9 Changed 10 years ago by emca

  • Description modified (diff)
  • Other processors set to emca

Eagle, Hawk and Camera data was delivered by anch 11/07/2011

Starting LiDAR processing

comment:10 Changed 10 years ago by emca

LiDAR data ready for delivery check

comment:11 Changed 10 years ago by adbe

LiDAR data ready to deliver

comment:12 Changed 10 years ago by emca

  • Description modified (diff)

LiDAR data has been dispatched 22/09/2011.
ASCII navigation information has also been included with the delivery as per PI's request.

comment:13 Changed 10 years ago by mark1

Hyperspectral hawk bad pixel files and XML files have been re-created and re-delivered to PI (via email). This corrects sample number in bad pixel text files and Eagle IFOV value in XML and extra comments included.

comment:14 Changed 10 years ago by adbe

Hyperspectral sct's

lineeaglehawk
1-0.01-0.05
2-0.05-0.05
3-0.05-0.06
4-0.02-0.06
5-0.03-0.07
6-0.02-0.04
7-0.02-0.05
8-0.01-0.05
9-0.08-0.05
10-0.02-0.05
11-0.04-0.05
12-0.03-0.06
13-0.03-0.06
14-0.05-0.05
15-0.05-0.05
16-0.031.06
17-0.06-0.05
18-0.05-0.05
19-0.04-0.05
20-0.03-0.05
21-0.02-0.06
22-0.04-0.05

comment:15 Changed 10 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from iopa to knpa

Preparing for archiving.

comment:16 Changed 10 years ago by knpa

No lidar intensity screenshot with vectors.

comment:17 Changed 10 years ago by knpa

Finished preparing for archiving.

comment:18 Changed 10 years ago by knpa

Backed-up by NEODC, moved to archive disk.

comment:19 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed
Note: See TracTickets for help on using tickets.